OpenStudy (anonymous):

Can someone help me walk through the steps of this problem? to help with future ones?

OpenStudy (anonymous):

Evaluate \[\sum_{n=1}^{4} 3n+1\] a) 13 b) 34 c) 3640 d) 1/13

OpenStudy (anonymous):

It's A

OpenStudy (anonymous):

I lied it's 34.. B

OpenStudy (unklerhaukus):

\[\sum_{n=1}^{4} 3n+1=\big[3(1)+1\big]+\big[3(2)+1\big]+\big[3(3)+1\big]+\big[3(4)+1\big]\\\qquad\quad\qquad=\]

OpenStudy (unklerhaukus):

a simpler example \[\sum_{n=4}^6n+1 =[4+1]+[5+1]+[6+1]=5+6+7=18\]

OpenStudy (anonymous):

oh, okay thank you for showing me the steps to it now it makes much more sense to how to look and solve a problem like this. Thank you

OpenStudy (anonymous):

so 34, thank you

OpenStudy (unklerhaukus):

yes!

OpenStudy (unklerhaukus):

The general sum \[\sum_{n=a}^Nf(n) = f(a)+f(a+1)+f(a+2)+\cdots+f(N-1)+f(N)\]